LASER INDUCED FLUORESCENCE OF PHTHALOCYANINE IN HYDROGEN DROPLETS

\underline{SUSUMU KUMA}, TAKAMASA MOMOSE, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, British Columbia V6T 1Z3, Canada..

Droplets of hydrogen molecules were generated by the same method as that of He droplets^{a}. Hydrogen droplets thus produced consist of about 10^{5} molecules with a temperature of about 4 K^{b}. We found that hydrogen droplets pick up phthalocyanine molecules in a pickup chamber like He droplets. Laser induced fluorescence spectra of phthalocyanine molecules captured by hydrogen droplets showed spectra similar to those observed for phthalocyanine-(H_{2})_{N} clusters embedded in He droplets with N > 1000, but the spectrum in hydrogen droplets showed sharper linewidth than that of phthalocyanine-(H_{2})_{N} in He droplets. In addition, a clear dependence in the linewidth on the ortho-to-para concentration ratio was observed. We will discuss details of the observed spectra.

aJ. P. Toennies and A. F. Vilesov, Angew. Chem. Int. Ed. 43, 2622 (2004)

bE. L. Knuth, F. Schünemann and J. P. Toennies, J. Chem. Phys. 102, 6258 (1995)